# Right on Time: Revising Time Series Models by Constraining their Explanations

### v3

@@ 4.2 Evaluations
Figure 4: Applying RioT makes the model ignore annotated confounder areas. While FCN primarily focuses on confounder areas, applying RioT with partial feedback (middle) or full feedback (bottom) causes the model to ignore the confounder and focus on the remainder of the input.
Removing Confounders for Time Series Classification. We evaluate the effectiveness of RioT (spatial: RioTsp, frequency: RioTfreq) in addressing confounders in classification tasks by comparing balanced accuracy with and without RioT. As shown in Tab. 1, without RioT, both FCN and OFA overfit to shortcuts, achieving $\approx$100training accuracy, while having poor test performance. Applying RioT significantly improves test performance for both models across all datasets. In some cases, RioT even reaches the performance of the ideal reference (unconfounded) scenario as if there would be no confounder in the data. Even on FordB, where the drop in training-to-test performance of the reference indicates a distribution shift, RioTsp is still beneficial. Similarly, RioTfreq enhances performance on frequency-confounded data, though the improvement is less pronounced for FCN on Ford B, suggesting essential frequency information is sometimes obscured by RioTfreq. In summary, RioT (both RioTsp and RioTfreq) successfully mitigates confounders, enhancing test generalization for FCN and OFA models.
@@
SP Conf 0.24 ±0.01 0.55 ±0.01 0.33 ±0.03 0.94 ±0.02 0.09 ±0.01 0.16 ±0.04
+ RioTsp 0.30 ±0.01 $\bullet$ 0.50 ±0.01 0.45 ±0.03 $\bullet$ 0.58 ±0.01 0.11 ±0.01 $\bullet$ 0.09 ±0.02
Freq Conf 0.30 ±0.02 0.46 ±0.01 0.33 ±0.04 0.36 ±0.04 0.11 ±0.02 0.32 ±0.09
+ RioTfreq 0.31 ±0.02 $\bullet$ 0.45 ±0.01 0.33 ±0.04 $\bullet$ 0.34 ±0.04 0.81 ±0.48 $\bullet$ 0.17 ±0.01
@@
SP Conf 0.40 ±0.02 0.55 ±0.01 0.29 ±0.01 0.96 ±0.03 0.20 ±0.03 0.19 ±0.01
+ RioTsp 0.40 ±0.03 $\bullet$ 0.53 ±0.01 0.44 ±0.00 $\bullet$ 0.45 ±0.01 0.55 ±0.20 $\bullet$ 0.14 ±0.01
Freq Conf 0.45 ±0.03 0.91 ±0.16 0.04 ±0.00 0.53 ±0.05 0.63 ±0.09 0.24 ±0.04
+ RioTfreq 0.91 ±0.07 $\bullet$ 0.66 ±0.04 2.45 ±4.59 $\bullet$ 0.38 ±0.06 0.96 ±0.02 $\bullet$ 0.17 ±0.00
@@
SP Conf 0.22 ±0.01 0.54 ±0.03 0.28 ±0.01 1.19 ±0.03 0.22 ±0.03 0.15 ±0.01
+ RioTsp 0.23 ±0.01 $\bullet$ 0.48 ±0.01 0.53 ±0.02 $\bullet$ 0.52 ±0.02 0.25 ±0.03 $\bullet$ 0.11 ±0.01
Freq Conf 0.06 ±0.01 0.69 ±0.08 0.07 ±0.01 0.34 ±0.08 0.79 ±0.09 0.31 ±0.09
+ RioTfreq 0.07 ±0.01 $\bullet$ 0.49 ±0.07 0.07 ±0.01 $\bullet$ 0.21 ±0.02 1.12 ±0.36 $\bullet$ 0.22 ±0.01
Removing Confounders in the Real-World. So far, our experiments have demonstrated the ability to counteract confounders within controlled environments. However, real-world scenarios often have more complex confounder structures. Our new proposed dataset P2S presents such real-world conditions. The model explanations for a sample in Fig. 4 (top) reveal a focus on distinct regions of the sensor curve, specifically the two middle regions. With domain knowledge, it’s clear that these regions shouldn’t affect the model’s output. By applying RioT, we can redirect the model’s attention away from these regions. New model explanations highlight that the model still focuses on incorrect regions, which can be mitigated by extending the annotated area. In Tab. 4, the model performance (exemplarly with FCN) in these settings is presented. Without RioT, the model overfits to the confounder. the test performance improves already with partial feedback (2) and improves even more with full feedback (4). These results highlight the effectiveness of RioT in real-world scenarios, where not all confounders are initially known.
@@
FCNsp + RioTsp (4) 0.95 ±0.01 $\bullet$ 0.82 ±0.06
Table 4: RioT can combine spatial and frequency feedback. If the data is confounded in the time and frequency domain, RioT can combine feedback on both domains to mitigate confounders, superior to feedback on only one domain. Unconfounded represents the ideal scenario when the model is not affected by any confounder.
@@
TiDEfreq,sp + RioTfreq,sp 0.25 ±0.01 $\bullet$ 0.47 ±0.01
We thus investigate the impact of applying RioT to both spatial and frequency confounders simultaneously (cf. Tab. 4), exemplary using FCN and TiDE. When Sleep is confounded in both domains, FCN without RioT overfits and fails to generalize. Addressing only one confounder does not mitigate the effects, as the model adapts to the other. However, combining feedback for both domains (RioTfreq,sp) significantly improves test performance, matching the frequency-confounded scenario (cf. Tab. 1). Tab. 4 (bottom) shows the impact of multiple confounders on the Energy dataset for forecasting. When faced with both spatial shortcut and noise confounders, the model overfits, indicated by lower training MSE. While applying either spatial or frequency feedback individually already shows some effect, utilizing both types of feedback simultaneously (RioTfreq,sp) results in the largest improvements, as both confounders are addressed. The performance gap between RioTfreq,sp and the non-confounded model is more pronounced than in single confounder cases (cf. Tab. 2), suggesting a compounded challenge. Optimize the deconfounding process in highly complex data environments thus remains an important challenge.
@@
Feedback Generalization. As human feedback is an essential aspect of RioT, we investigate the required annotations and the potential to generalize annotations across samples. Our findings indicate that not every sample needs annotation. Fig. 5 shows that we can significantly reduce the amount of annotated data for classification and forecasting (cf. App. Tab. 6 and Tab. 5 for results on the other datasets). Even minimal feedback, such as annotating just 5% of the samples, substantially improves performance compared to no feedback. Furthermore, the results on P2S highlights that annotations can be generalized across multiple samples. Once the confounder on P2S has been identified on a couple of samples, the expert annotations can be used on full dataset. The systematic nature of shortcut confounders suggest that generalizing annotations is an effective possibility to obtain feedback efficiently. While RioT does rely on human annotations, these findings highlight that it can work without extensive manual human interactions, and that obtained annotations can be utilized efficiently.
@@ 5 Conclusion
In this work, we present Right on Time a method to mitigate confounding factors in time series data with the help of human feedback. By revising the model, RioT significantly diminishes the influence of these factors, steering the model to align with the correct reasons. Using popular time series models on several manually confounded datasets and the newly introduced, naturally confounded, real-world dataset P2S showcases that they are indeed subject to confounders. Our results, however, demonstrate that applying RioT to these models can mitigate confounders in the data. Furthermore, we have unveiled that addressing solely the time domain is insufficient for revising the model to focus on the correct reasons, which is why we extended our method beyond it. Feedback in the frequency domain provides an additional way to steer the model away from confounding factors and towards the right reasons. Extending the application of RioT to multivariate time series represents a logical next step, and exploring the integration of various explainer types is another promising direction. Additionally, we aim to apply RioT, especially RioTfreq, to other modalities as well, offering a more nuanced approach to confounder mitigation. It should be noted that while our method shows potential in its current iteration, interpreting attributions in time series data remains a general challenge.
@@ Appendix A Appendix
A.1 Impact Statement
Our research advances machine learning by enhancing the interpretability and reliability of time series models, significantly impacting human interaction with AI systems. By developing Right on Time (RioT), which guides models to focus on correct reasoning, we improve the transparency and trust in machine learning decisions. While human feedback can provide many benefits, one has also to be aware that it could be incorrect, and evaluate the consequences carefully.
A.2 Implementation and Experimental Details
Adaption of Integrated Gradients (IG). A part of IG is a multiplication of the model gradient with the input itself, improving the explanation’s quality [Shrikumar et al., 2017]. However, this multiplication makes some implicit assumptions about the input format. In particular, it assumes that there are no inputs with negative values. Otherwise, multiplying the attribution score with a negative input would flip the attribution’s sign, which is not desired. For images, this is unproblematic because they are always equal to or larger than zero. In time series, negative values can occur and normalization to make them all positive is not always suitable. To avoid this problem, we use only the input magnitude and not the input sign to compute the IG attributions.
@@ A.2 Implementation and Experimental Details
Model Training and Hyperparameters. To find suitable parameters for model training, we performed a hyperparameter search over batch size, learning rate, and the number of training epochs. We then used these parameters for all model trainings and evaluations, with and without RioT. In addition, we selected suitable $\lambda$ values for RioT with a hyperparameter selection on the respective validation sets. The exact values for the model training parameters and the $\lambda$ values can be found in the provided code.
@@ A.3 UCR Dataset selection
A.4 Computational Costs of RioT
@@
where $g=\frac{\partial\mathcal{L}_{\mathrm{RA}}}{\partial\theta}$ is the partial derivative of the right answer loss and if $H$ is the full joint Hessian matrix of the loss with respect to $\theta$ and $x$, then $H_{\theta x}$ is the sub-block of this matrix mapping $x$ into $\theta$ (cf. Fig. 6), given by $H_{\theta x}=\frac{\partial^{2}f_{\theta}(x)}{\partial\theta\partial x}$. Hessian-vector products are known to be fast to compute [Martens, 2010], enabling the right-reason loss computation to scale to large models and inputs.
@@
A.5 Details on Confounding Factors
@@
while $t\in\{0,1,\dots,T\}$ and $j$ represents the class index, simulating a spurious correlation between the confounder and class index.
@@
Forecasting Spatial. For forecasting datasets, spatial confounders are shortcuts that act as the actual solution to the forecasting problem. Periodically, data from the time series is copied back in time. This “back-copy” is a shortcut for the forecast, as it resembles the time steps of the forecasting window. Due to the windowed sampling from the time series, this shortcut occurs at every second sample. The exact confounder formulation is outlined in the sketch below (Fig. 7), with an exemplary lookback length of $9$, forecasting horizon of $3$ and window stride of $6$. This results in a shortcut confounder in samples 1 and 3 (marked red) and overlapping in sample 2 (marked orange).
@@
A.6 Additional Experimental Results
@@
SP Conf 0.34 ±0.01 0.54 ±0.01 0.44 ±0.03 0.77 ±0.01 0.21 ±0.01 0.30 ±0.04
+ RioTsp 0.40 ±0.01 $\bullet$ 0.52 ±0.01 0.53 ±0.02 $\bullet$ 0.62 ±0.01 0.23 ±0.01 $\bullet$ 0.22 ±0.01
Freq Conf 0.39 ±0.01 0.47 ±0.01 0.45 ±0.03 0.45 ±0.03 0.21 ±0.03 0.45 ±0.06
+ RioTfreq 0.40 ±0.01 $\bullet$ 0.47 ±0.01 0.45 ±0.03 $\bullet$ 0.44 ±0.02 0.59 ±0.22 $\bullet$ 0.39 ±0.01
@@
SP Conf 0.46 ±0.00 0.53 ±0.01 0.41 ±0.00 0.78 ±0.01 0.32 ±0.04 0.33 ±0.00
+ RioTsp 0.46 ±0.01 $\bullet$ 0.52 ±0.01 0.51 ±0.00 $\bullet$ 0.53 ±0.01 0.54 ±0.12 $\bullet$ 0.28 ±0.00
Freq Conf 0.53 ±0.01 0.81 ±0.07 0.15 ±0.00 0.64 ±0.03 0.58 ±0.03 0.41 ±0.05
+ RioTfreq 0.92 ±0.05 $\bullet$ 0.80 ±0.02 0.97 ±0.86 $\bullet$ 0.57 ±0.02 0.65 ±0.01 $\bullet$ 0.40 ±0.01
@@
SP Conf 0.32 ±0.01 0.54 ±0.01 0.40 ±0.01 0.85 ±0.01 0.32 ±0.03 0.29 ±0.01
+ RioTsp 0.34 ±0.01 $\bullet$ 0.51 ±0.01 0.57 ±0.01 $\bullet$ 0.58 ±0.01 0.35 ±0.03 $\bullet$ 0.24 ±0.01
Freq Conf 0.18 ±0.01 0.74 ±0.06 0.18 ±0.01 0.53 ±0.07 0.65 ±0.05 0.49 ±0.09
+ RioTfreq 0.19 ±0.01 $\bullet$ 0.60 ±0.05 0.18 ±0.01 $\bullet$ 0.40 ±0.03 0.79 ±0.16 $\bullet$ 0.41 ±0.02
@@
TiDEfreq,sp + RioTfreq,sp 0.39 ±0.01 $\bullet$ 0.55 ±0.01
This section provides further insights into our experiments, covering both forecasting and classification tasks. Specifically, it showcases performance through various metrics such as MAE, MSE, and accuracy, and explores different feedback configurations.

### v4

@@ 4.2 Evaluations
Removing Confounders for Time Series Classification.
Figure 4: Applying RioT lets the model ignore confounder areas. While FCN primarily focuses on confounders, applying RioT with partial feedback (middle) or full feedback (bottom) causes the model to ignore the confounder and focus on the remainder of the input.
We evaluate the effectiveness of RioT (spatial: RioTsp, frequency: RioTfreq) in addressing confounders in classification tasks by comparing balanced accuracy with and without RioT.
As shown in Tab. 1, without RioT, both FCN and OFA overfit to shortcuts, achieving $\approx$100training accuracy, while having poor test performance. Applying RioT significantly improves test performance for both models across all datasets. In some cases, RioT even reaches the performance of the ideal reference (unconfounded) scenario as if there would be no confounder in the data. Even on FordB, where the drop in training-to-test performance of the reference indicates a distribution shift, RioTsp is still beneficial. Similarly, RioTfreq enhances performance on frequency-confounded data, though the improvement is less pronounced for FCN on Ford B, suggesting essential frequency information is sometimes obscured by RioTfreq. In summary, RioT (both RioTsp and RioTfreq) successfully mitigates confounders, enhancing test generalization for FCN and OFA models.
@@
Basesp 0.24 ±0.01 0.55 ±0.01 0.33 ±0.03 0.94 ±0.02 0.09 ±0.01 0.16 ±0.04
+ RioTsp 0.30 ±0.01 0.50 ±0.01 0.45 ±0.03 0.58 ±0.01 0.11 ±0.01 0.09 ±0.02
Basefreq 0.30 ±0.02 0.46 ±0.01 0.33 ±0.04 0.36 ±0.04 0.11 ±0.02 0.32 ±0.09
+ RioTfreq 0.31 ±0.02 0.45 ±0.01 0.33 ±0.04 0.34 ±0.04 0.81 ±0.48 0.17 ±0.01
@@
Basesp 0.40 ±0.02 0.55 ±0.01 0.29 ±0.01 0.96 ±0.03 0.20 ±0.03 0.19 ±0.01
+ RioTsp 0.40 ±0.03 0.53 ±0.01 0.44 ±0.00 0.45 ±0.01 0.55 ±0.20 0.14 ±0.01
Basefreq 0.45 ±0.03 0.91 ±0.16 0.04 ±0.00 0.53 ±0.05 0.63 ±0.09 0.24 ±0.04
+ RioTfreq 0.91 ±0.07 0.66 ±0.04 0.71 ±0.10 0.38 ±0.06 0.96 ±0.02 0.17 ±0.00
@@
Basesp 0.22 ±0.01 0.54 ±0.03 0.28 ±0.01 1.19 ±0.03 0.22 ±0.03 0.15 ±0.01
+ RioTsp 0.23 ±0.01 0.48 ±0.01 0.53 ±0.02 0.52 ±0.02 0.25 ±0.03 0.11 ±0.01
Basefreq 0.06 ±0.01 0.69 ±0.08 0.07 ±0.01 0.34 ±0.08 0.79 ±0.09 0.31 ±0.09
+ RioTfreq 0.07 ±0.01 0.49 ±0.07 0.07 ±0.01 0.21 ±0.02 1.12 ±0.36 0.22 ±0.01
Removing Confounders in the Real-World. So far, our experiments have demonstrated the ability to counteract confounders within controlled environments. However, real-world scenarios often have more complex confounder structures. Our new proposed dataset P2S presents such real-world conditions. The model explanations in Fig. 4 (top) reveal a focus on distinct regions of the sensor curve, specifically the two middle regions. With domain knowledge, it’s clear that these regions shouldn’t affect the model’s output. By applying RioT, we can redirect the model’s attention away from these regions. New model explanations highlight that the model still focuses on incorrect regions, which can be mitigated by extending the annotated area. In Tab. 4, the model performance (exemplarly with FCN) in these settings is presented. Without RioT, the model overfits to the confounder. the test performance improves already with partial feedback (2) and improves even more with full feedback (4). These results highlight the effectiveness of RioT in real-world scenarios, where not all confounders are initially known.
@@
FCNsp + RioTsp (4) 0.95 ±0.01 0.82 ±0.06
Table 4: RioT can combine spatial and frequency feedback. If the data is confounded in time and frequency, RioT can combine feedback on both domains to mitigate confounders, superior to feedback on only one domain. Unconfounded represents the ideal scenario when the model is not affected by any confounder.
@@
TiDEfreq,sp + RioTfreq,sp 0.25 ±0.01 0.47 ±0.01
We thus investigate the impact of applying RioT to both spatial and frequency confounders simultaneously (cf. Tab. 4), exemplary using FCN and TiDE. When Sleep is confounded in both domains, FCN without RioT overfits and fails to generalize. Addressing only one confounder does not mitigate the effects, as the model adapts to the other. However, combining the respective feedback from both domains (RioTfreq,sp) significantly improves test performance, matching the frequency-confounded scenario (cf. Tab. 1). Tab. 4 (bottom) shows the impact of multiple confounders on the Energy dataset for forecasting. When faced with both spatial shortcut and noise confounders, the model overfits, indicated by lower training MSE. While applying either spatial or frequency feedback individually already shows some effect, utilizing both types of feedback simultaneously (RioTfreq,sp) results in the largest improvements, as both confounders are addressed. The performance gap between RioTfreq,sp and the non-confounded model is more pronounced than in single confounder cases (cf. Tab. 2), suggesting a compounded challenge. Optimize the deconfounding process in highly complex data environments thus remains an important challenge.
@@
Figure 6: RioT is robust against invalid feedback. Even with some percentage of random feedback, RioT overcomes the confounders.
Handling Human Feedback. Human feedback is a crucial component of RioT. To understand its impact, we conduct two ablation studies using the classification data set Fault Detection A and the forecasting data set Energy. The first experiment examines the required amount of feedback, while the second assesses robustness to noisy feedback.
Recognizing that expert time is valuable and excessive feedback requests are impractical, our first experiment evaluates RioT ’s performance when feedback is provided on only a portion of the dataset (Fig. 6). The findings reveal that full annotations are unnecessary. Even with minimal feedback, such as annotating just 5% of the samples, RioT significantly outperforms scenarios with no feedback.
While previous experiments assumed entirely accurate feedback, real-world applications often involve some degree of error. Therefore, we also test RioT ’s resilience to increasing levels of incorrect feedback (Fig. 6). Instead of accurately marking confounding areas, random time steps or frequency components are incorrectly labeled as confounded. The results show that RioT maintains strong performance even with up to 10% invalid feedback, presenting only slight performance declines. In certain cases, like forecasting with spatial confounders, RioT can still achieve notable improvements despite high levels of feedback noise.
In summary, RioT effectively generalizes from small subsets of feedback and remains robust against a moderate amount of annotation noise. These qualities demonstrate that RioT is well-equipped to manage the practical challenges associated with human feedback.
@@ 5 Conclusion
In this work, we present Right on Time a method to mitigate confounding factors in time series data with the help of human feedback. By revising the model, RioT significantly diminishes the influence of these factors, steering the model to align with the correct reasons. Using popular time series models on several manually confounded datasets and the newly introduced, naturally confounded, real-world dataset P2S showcases that they are indeed subject to confounders. Our results, however, demonstrate that applying RioT to these models can mitigate confounders in the data. Furthermore, we have unveiled that addressing solely the time domain is insufficient for revising the model to focus on the correct reasons, which is why we extended our method beyond it. Feedback in the frequency domain provides an additional way to steer the model away from confounding factors and towards the right reasons. Extending the application of RioT to multivariate time series represents a logical next step, and exploring the integration of various explainer types is another promising direction. To further reduce the required human annotations, exploring the use of semi-automated feeback techniques, transfer learning or LLMs are potential next steps. Additionally, we aim to apply RioT, especially RioTfreq, to other modalities as well, offering a more nuanced approach to confounder mitigation. It should be noted that while our method shows potential in its current iteration, interpreting attributions in time series data remains a general challenge.
6 Ethics Statement
Our research aims to enhance the interpretability and reliability of time series models, with a focus on improving human interaction with time series models. By developing RioT, we prioritize guiding models toward correct reasoning, increasing transparency and trust in machine learning decisions. While human feedback plays a crucial role in refining these models, we acknowledge the potential for inaccuracies in the feedback. In production settings, it is crucial to consider these potential risks and implement appropriate safeguards to ensure responsible and reliable AI deployment.
7 Reproducibility Statement
To ensure the reproducibility of our research, we have made the code [^3] https://github.com/ml-research/RioTand dataset  [^4] https://huggingface.co/datasets/AIML-TUDA/P2S used in this work publicly available. Detailed instructions for running the experiments and replicating the results are provided in the repository, along with any dependencies and configurations required. The methodology is thoroughly described in the appendix SubSec. A.1, outlining the steps implement and evaluate RioT. We encourage others to use these resources to validate and extend our findings.
@@ Appendix A Appendix
A.1 Implementation and Experimental Details
Adaption of Integrated Gradients (IG). A part of IG is a multiplication of the model gradient with the input itself, improving the explanation’s quality (Shrikumar et al., 2017). However, this multiplication makes some implicit assumptions about the input format. In particular, it assumes that there are no inputs with negative values. Otherwise, multiplying the attribution score with a negative input would flip the attribution’s sign, which is not desired. For images, this is unproblematic because they are always equal to or larger than zero. In time series, negative values can occur and normalization to make them all positive is not always suitable. To avoid this problem, we use only the input magnitude and not the input sign to compute the IG attributions.
@@ A.1 Implementation and Experimental Details
Model Training and Hyperparameters. To find suitable parameters for model training, we performed a hyperparameter search over batch size, learning rate, and the number of training epochs. We then used these parameters for all model trainings and evaluations, with and without RioT. In addition, we selected suitable $\lambda$ values for RioT with a hyperparameter selection on the respective validation sets. The exact values for the model training parameters and the $\lambda$ values can be found in the provided code [^6] https://github.com/ml-research/RioT.
@@ A.2 UCR Dataset selection
A.3 Computational Costs of RioT
@@
where $g=\frac{\partial\mathcal{L}_{\mathrm{RA}}}{\partial\theta}$ is the partial derivative of the right answer loss and if $H$ is the full joint Hessian matrix of the loss with respect to $\theta$ and $x$, then $H_{\theta x}$ is the sub-block of this matrix mapping $x$ into $\theta$ (cf. Fig. 7), given by $H_{\theta x}=\frac{\partial^{2}f_{\theta}(x)}{\partial\theta\partial x}$. Hessian-vector products are known to be fast to compute (Martens, 2010), enabling the right-reason loss computation to scale to large models and inputs.
@@
A.4 Details on Confounding Factors
@@
while $t\in\{0,1,\dots,T\}$ and $j$ represents the class index, simulating a spurious correlation between the confounder and class index (Fig. 8).
Figure 8: Example of the added spatial confounder in the Fault Detection dataset.
@@
Figure 9: Example of an added frequency shortcut in the Fault Detection dataset.
Forecasting Spatial. For forecasting datasets, spatial confounders are shortcuts that act as the actual solution to the forecasting problem. Periodically, data from the time series is copied back in time. This “back-copy” is a shortcut for the forecast, as it resembles the time steps of the forecasting window. Due to the windowed sampling from the time series, this shortcut occurs at every second sample. The exact confounder formulation is outlined in the sketch below (Fig. 10), with an exemplary lookback length of $9$, forecasting horizon of $3$ and window stride of $6$. This results in a shortcut confounder in samples 1 and 3 (marked red) and overlapping in sample 2 (marked orange). This can for example occur in scenarios such as data transmission. Transmission glitches, manifesting as packet losses or duplications, can subtly introduce irregularities into the time series data.
@@
Such a confounder could, for example, occur when monitoring of water flow through a pipe in an assembly line. Suppose that during the capturing, a defect in one of the rollers adjacent to the water pipe induces a systematic stutter. This mechanical stutter, in turn, generates repeated impulses in the water flow sensor’s readings. These impulses act as a systematic frequency confounder, which can negatively influence the performance of the forecasting model.
@@
A.5 Additional Experimental Results
Figure 11: t-SNE plots of OFA encodings for Fault Detection A. The left plot shows a confounded model with minimal class separation. The middle plot shows a confounded model after RioT regularization, while the far right plot shows an unconfounded model with clear class separation. Both RioT-regularized and unconfounded model exhibit similar structures, highlighting the effectiveness of RioT.
@@ A.5 Additional Experimental Results
Basesp 0.34 ±0.01 0.54 ±0.01 0.44 ±0.03 0.77 ±0.01 0.21 ±0.01 0.30 ±0.04
+ RioTsp 0.40 ±0.01 0.52 ±0.01 0.53 ±0.02 0.62 ±0.01 0.23 ±0.01 0.22 ±0.01
Basefreq 0.39 ±0.01 0.47 ±0.01 0.45 ±0.03 0.45 ±0.03 0.21 ±0.03 0.45 ±0.06
+ RioTfreq 0.40 ±0.01 0.47 ±0.01 0.45 ±0.03 0.44 ±0.02 0.59 ±0.22 0.39 ±0.01
@@
Basesp 0.46 ±0.00 0.53 ±0.01 0.41 ±0.00 0.78 ±0.01 0.32 ±0.04 0.33 ±0.00
+ RioTsp 0.46 ±0.01 0.52 ±0.01 0.51 ±0.00 0.53 ±0.01 0.54 ±0.12 0.28 ±0.00
Basefreq 0.53 ±0.01 0.81 ±0.07 0.15 ±0.00 0.64 ±0.03 0.58 ±0.03 0.41 ±0.05
+ RioTfreq 0.92 ±0.05 0.80 ±0.02 0.97 ±0.86 0.57 ±0.02 0.65 ±0.01 0.40 ±0.01
@@
Basesp 0.32 ±0.01 0.54 ±0.01 0.40 ±0.01 0.85 ±0.01 0.32 ±0.03 0.29 ±0.01
+ RioTsp 0.34 ±0.01 0.51 ±0.01 0.57 ±0.01 0.58 ±0.01 0.35 ±0.03 0.24 ±0.01
Basefreq 0.18 ±0.01 0.74 ±0.06 0.18 ±0.01 0.53 ±0.07 0.65 ±0.05 0.49 ±0.09
+ RioTfreq 0.19 ±0.01 0.60 ±0.05 0.18 ±0.01 0.40 ±0.03 0.79 ±0.16 0.41 ±0.02
@@
TiDEfreq,sp + RioTfreq,sp 0.39 ±0.01 0.55 ±0.01
This section provides further insights into our experiments, covering both forecasting and classification tasks. Specifically, it showcases performance through various metrics such as MAE, MSE, and accuracy, qualitative insights about the influence of confounders, and explores different feedback configurations.
Qualitative Insights into Encodings of a Confounded Model. In Fig. 11, t-SNE plots outline the feature encodings of OFA trained on Fault Detection A under three different scenarios. The plot on the left represents a model trained on confounded data without any regularization, showing minimal class separation and no discernible structure, indicating poor feature representation caused by confounding factors. The middle plot depicts the same confounded model after applying RioT regularization, where class separation and improved structure emerge. The far right plot displays the eoncodings of a model trained on unconfounded data, with clear and distinct class clusters.
The qualitative insights are further supported by the scores presented in Tab. 1, which detail the corresponding classification performance of the presented models. The OFA model trained on confounded data achieves only $\approx$50accuracy, while the RioT-regularized model regains nearly 100% accuracy, comparable to the unconfounded model. This improvement in accuracy aligns with the latent representations observed in the t-SNE plots, where RioT effectively steers the confounded model’s structure to resemble that of the unconfounded data. These results highlight RioT’s capability to mitigate confounding and restore robust model performance.
@@
Early Stopping as Confounder Mitigation Baseline. In this experiment, we compare the performance of a model with RioT to a model regularized via early stopping (which is decided on an unconfounded validation set). In that, we stop model training if there are no improvements in the validation set for several epochs in the hope that it thus does not overfit to the confounder. The results are presented in Tab. 9 for classification and Tab. 10 for forecasting. We can observe that early stopping can help in some instances to achieve performances similar to RioT (e.g. PatchTST with a frequency confounder or FCN with a spatial confounder). However, for the majority of cases the performance with early stopping is substantially lower than the performance with RioT, signaling that early stopping alone is not a sufficient approach to overcome confounders.
Config FCN OFA
Train Test Train Test
Unconfounded $0.99\pm 0.00$ $0.99\pm 0.00$ $1.00\pm 0.00$ $0.98\pm 0.02$
Basesp $1.00\pm 0.00$ $0.74\pm 0.06$ $1.00\pm 0.00$ $0.53\pm 0.02$
[…]
Basefreq $0.98\pm 0.01$ $0.87\pm 0.03$ $1.00\pm 0.00$ $0.72\pm 0.02$
[…]
Table 9: Early stopping on classification datasets.
Config PatchTST TiDE
Train Test Train Test
Unconfounded $0.26\pm 0.01$ $0.23\pm 0.00$ $0.27\pm 0.01$ $0.26\pm 0.02$
Basesp $0.29\pm 0.01$ $0.96\pm 0.03$ $0.28\pm 0.01$ $1.19\pm 0.03$
[…]
Basefreq $0.04\pm 0.00$ $0.53\pm 0.05$ $0.07\pm 0.01$ $0.34\pm 0.08$
[…]
Table 10: Early stopping on forecasting datasets.